# Geometric Height on flag varieties in positive characteristic

### v1

@@ Abstract.
Let $k$ be an algebraically closed field of characteristic $p\neq 0$. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow G$ be a strictly anti-dominant character. Let $C$ be a projective smooth curve over $k$ with function field $K=k(C)$ and $F$ be a principal $G$-bundle on $C$. Then $F/P\longrightarrow C$ is a flag bundle and $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$ is a relatively ample line bundle. We compute the height filtration and successive minima of the height function $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ over the flag variety $X=(F/P)_{K}$.
@@ 1.1. Height filtration and successive minima
Let $K$ be either a number field or $K=k(C)$ where $C$ is a projective smooth curve over a field $k$. Let $X$ be a projective variety of dimension $d$ over $K$ and $\overline{L}$ be an adelic line bundle on $X$. These data induce an Arakelov height function $h_{\overline{L}}$ on $X$ (see [9, §9] for a survey). A typical case is the geometric height, which is the one we concern in this article. Here we give the definition.
If $K=k(C)$, consider a projective flat morphism $\mathcal{X}\longrightarrow C$ with the generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and a line bundle $\mathcal{L}$ on $\mathcal{X}$ with $\mathcal{L}_{K}\simeq L$. The data $(\mathcal{X},\mathcal{L})$ define an adelic line bundle $\overline{L}$ and the height function $h_{\overline{L}}$ is given by
$h_{\overline{L}}:X(\overline{K})\longrightarrow\mathbb{Q},\;x\longmapsto\frac{\mathcal{L}\cdot\overline{\{x\}}}{{\mathrm{deg}}(x)}\;\;\text{where $\overline{\{x\}}$ is the closure of $x$ in $\mathcal{X}$}.$
We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. If $K$ is a number field, the height function can be defined similarly via arithmetic intersection theory.
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big{\{}x\in X(\overline{K}):h_{\overline{L}}(x)<t\big{\}}$. Let’s call $\big{\{}Z_{t}:t\in\mathbb{R}\big{\}}$ the height filtration, and call its jumping points the successive minima. Note that our definition of successive minima are slightly different with Zhang [10]. Zhang considers only dimension jumps $e_{i}=\inf\big{\{}t:\dim Z_{t}\geq d-i+1\big{\}}$, $i=1,\dots,d+1$. In this article, $e_{i}$ will be called the successive minima of Zhang to avoid ambiguity.
In [3], Fan-Luo-Qu provides a new case where height filtration can be explicitly computed, which is the geometric height on flag varieties under the assumption ${\rm char}(k)=0$. Following their essence, we examine the case of positive characteristic.
Let $k$ be an algebraically closed field and $C$ be a projective smooth curve over $k$ with function field $K=k(C)$. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $F$ be a principal $G$-bundle over $C$ and $\mathcal{X}=F/P$ with generic fiber $X=(F/P)_{K}$. Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function by the following diagram.
@@
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. The numbers $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are well-defined. One main result in [3] is the following.
@@
From now on assume $k$ has positive characteristic unless other description is given. We show that this result does not hold in characteristic $p$ unless one puts the following assumption on $F$, namely $F$ admits a strongly canonical reduction.
@@
Canonical reduction is automatically strongly canonical in characteristic 00, whereas strongly canonical reduction may not exist in characteristic $p>0$. For general $G$-bundles that may not admit a strongly canonical reduction, we have the following rough form of our another main result.
@@
The height filtration of $X$ is given by successively deleting some Frobenius twist of Schubert cells.
More explicitly, a theorem by Langer [5] shows for a principal $G$-bundle $F$ the strongly canonical reduction exists for $({\rm Fr}^{n})^{*}F$, when $n$ is sufficiently large. Therefore, consider the following Cartesian diagram for $n$ sufficiently large.
${(\operatorname{Fr_{C}}^{n})^{*}F/P}$${F/P}$${C}$${C}$$\scriptstyle{\phi}$$\scriptstyle{\operatorname{Fr}_{C}^{n}}$
@@
[…]
Here $\operatorname{Fr}_{C}$ is the absolute Frobenius on $C$. Suppose $n$ is large enough such that $(\operatorname{Fr}_{C}^{n})^{*}F$ has strongly canonical reduction, then we have:
@@
The height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by the image of the height filtration of the height filtration of $\tilde{X}$ along the homeomorphism $\phi_{K}$, the successive minima of $X$ are $\frac{1}{p^{n}}$ of the succesive minima of $\tilde{X}$.
@@
Here the height filtration of $\tilde{X}$ is computed by Theorem 3.1. Thus, we say the height filtration of $X$ is given by successively deleting some Frobenius twist of Schubert cells as stated in Theorem 1.4.
@@ 1.2. Toy example: projective spaces
Let $k$ be an algebraically closed field of positive characteristic and $C$ be a curve over $k$ with function field $K=k(C)$. Let $E$ be a vector bundle of rank $n$ on $C$. We take $\mathcal{X}=\mathbb{P}(E)$, $\mathcal{L}$ to be the relative ample line bundle $\mathcal{O}_{\mathbb{P}(E)}(1)$ and $X=\mathbb{P}(V)$ where $V$ is the generic fiber of $E$. A special case of a theorem by Ballaÿ [1] show that: Let $\zeta_{ess}:=\inf_{t\in\mathbb{R}}\{Z_{t}=X\}$. Then
@@
where $\mu_{\max}(-)$ is the maximal slope appears in the Harder-Narasimhan filtration.
@@
In positive characteristic, the symmetric power of a semistable vector bundle may not remain semistable, making the left-hand side quite difficult to compute. It’s important to note that this issue arises only in positive characteristic. However, this problem can be resolved by sufficiently twisting the Frobenius morphism multiple times.
@@ 1.3. Acknowledgments
We sincerely thank Dr. Binggang Qu for his invaluable advice and support on this work. This research also stems from the 2024 Algebra and Number Theory Summer School held at Peking University. We are deeply grateful to the organizers for their efforts and hospitality.
@@ 2.1. Vector bundles
Let $E$ be a vector bundle on $C$. The degree of $E$ is ${\mathrm{deg}}(E):={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E):={\mathrm{deg}}(E)/\operatorname{rk}(E)$. It is called slope semistable if for every subbundle $F$ of $E$, $\mu(F)\leq\mu(E)$. This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
There exists uniquely a filtration $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
@@
This filtration is called the Harder-Narasimhan filtration of $E$. $\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$.
@@ 2.2. Principal bundles
A principal $\Gamma$-bundle on $C$ is a variety $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
@@
${\mathrm{deg}}(F):X(\Gamma)\longrightarrow\mathbb{Z},\quad\lambda\longmapsto\langle{\mathrm{deg}}(F),\lambda\rangle={\mathrm{deg}}(F\times_{\Gamma}k_{\lambda}).$
Here ${\mathrm{deg}}(F\times_{\Gamma}k_{\lambda})$ is the degree of the line bundle $F\times_{\Gamma}k_{\lambda}$ on the curve $C$.
Let $H$ be a closed subgroup of a linear algebraic group $\Gamma$ over $k$. A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\phi)$ where $F_{H}$ is a principal $H$-bundle and $\phi:F_{H}\times_{H}\Gamma\simeq F$ is an isomorphism.
By the universal property of the quotient $F/H$, the assignment to a section $\sigma:C\longrightarrow F/H$ the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between reductions of structure group of $F$ to $H$ and sections of $F/H\longrightarrow C$.
@@ 2.3. Reductive groups, characters and cocharacters
and by taking duals, we get the so-called slope map $X(P)^{\vee}_{\mathbb{Q}}\longrightarrow X(T)^{\vee}_{\mathbb{Q}}$ introduced in [8, §2.1.3]. In other words, a cocharacter on $X(P)_{\mathbb{Q}}$ can be extended canonically to $X(T)_{\mathbb{Q}}$.
2.4. The strongly canonical reduction to a parabolic subgroup
Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$. Let $F_{P}$ be a reduction of $F$ to a parabolic subgoup $P$. Let ${\mathrm{deg}}(F_{P})\in X(T)_{\mathbb{Q}}^{\vee}$ be the induced (rational) cocharacter.
The $G$-bundle $F$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$.
@@ 2.4. The strongly canonical reduction to a parabolic subgroup
Among all filtrations of a vector bundle, there is a canonical one (the Harder-Narasimhan filtration). Similarly, among all reduction to parabolic subgroups, there is a canonical one. A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
@@
We say $F_{Q}$ is strongly canonical if for any non-constant finite morphism $f:C^{\prime}\to C$, the pullback $f^{*}(F_{Q})$ is canonical.
@@
If the base field has characteristic 00, then semistability (resp. canonical reduction) and strongly semistability (resp. strongly canonical reduction) are equivalent and the (strongly) canonical reduction exists uniquely [2]. However, in positive characteristic, the notion of strongly canonical reduction becomes essential for our purposes, despite the fact that it does not generally exist.
@@ 3. Height filtrations and successive minima
${X=(F/P)_{K}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{X}=F/P\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Spek}(K)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C}$
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the degree cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. We have the following theorem.
@@
In particular, successive minima are $\zeta_{w}=\left\langle\operatorname{deg}\left(F_{Q}\right),w\lambda\right\rangle$ and Zhang’s successive minima are $e_{i}=\min\left\{\zeta_{w}:\ell(w)=\operatorname{dim}G/P-i+1\right\}$ where
@@
[…]

### v2

@@ Abstract.
Let $k$ be an algebraically closed field of characteristic $p\neq 0$. Let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly anti-dominant character. Let $C$ be a projective smooth curve over $k$ with function field $K=k(C)$ and $F$ be a principal $G$-bundle on $C$. We compute explicitly the height filtration and successive minima of the height function $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ over the flag variety $X=(F/P)_{K}$ through relatively ample line bundle the $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ on $F/P$.
@@ 1.1. Height filtration and successive minima
Let $K$ be either a number field or $K=k(C)$ where $C$ is a projective smooth curve over a field $k$. Let $X$ be a projective variety of dimension $d$ over $K$ and $\overline{L}$ be an adelic line bundle on $X$. These datum induce an Arakelov height function $h_{\overline{L}}$ on $X$ (see [11, §9] for a survey).
A typical case is the geometric height, which is the one we concern in this article. Here we give an equivalent definition.
In our case $K=k(C)$, suppose there is a projective flat morphism $\mathcal{X}\longrightarrow C$ with the generic fiber $X\longrightarrow\operatorname{Spec}(K)$ and a line bundle $\mathcal{L}$ on $\mathcal{X}$. The data $(\mathcal{X},\mathcal{L})$ define an adelic line bundle $\overline{L}$ and the height function $h_{\overline{L}}$ is given by
[…]
We also denote this by $h_{\mathcal{L}}$ if there is no ambiguity. For a different choice $(\mathcal{X},\mathcal{L})$, the height functions differ by a bounded function, see [11].
For any $t\in\mathbb{R}$, let $Z_{t}\subseteq X$ be the Zariski closure of the set $\big\{x\in X(\overline{K}):h_{\overline{L}}(x)<t\big\}$. We call $\big\{Z_{t}:t\in\mathbb{R}\big\}$ the height filtration, and call its jumping points the successive minima. Note that this definition of successive minima is slightly different from Zhang [12]. Zhang considers only dimension jumps $e_{i}=\inf\big\{t:\dim Z_{t}\geq d-i+1\big\}$, $i=1,\dots,d+1$.
In [4], Fan-Luo-Qu provides a new case where height filtration can be explicitly computed, which is the geometric height on flag varieties under the assumption ${\rm char}(k)=0$. Following their idea, we complete this calculation by examining this case in positive characteristic.
Following previous notations, let $G$ be a connected reductive group over $k$, $P\subseteq G$ be a parabolic subgroup, and $\lambda:P\longrightarrow\mathbb{G}_{m}$ be a strictly antidominant character. Let $F$ be a principal $G$-bundle over $C$ and $\mathcal{X}=F/P$ with generic fiber $X=(F/P)_{K}$. Let $\mathcal{L}_{\lambda}=F\times_{P}k_{\lambda}$ and $h_{\mathcal{L}_{\lambda}}:X(\overline{K})\longrightarrow\mathbb{R}$ be the induced height function by the following diagram.
@@
Denote $F_{Q}$ to be the canonical reduction of $F$, which is a principal $Q$-bundle for a parabolic subgroup $Q\subseteq G$, and also ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ to be the induced cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$ correspondingly. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell. The numbers $\langle{\mathrm{deg}}(F_{Q}),w\lambda\rangle$ are well-defined. One main result in [4] computes the height filtration in characteristic zero.
@@
From now on, we assume $k$ has positive characteristic unless other description is given. We compute the height filtration in this case. We show that this result does not hold in characteristic $p$ unless one puts the following assumption on $F$, namely $F$ admits a strongly canonical reduction.
@@
Canonical reduction is preseved by pullbacks along separated morphisms, which means the canonical reduction is strong in characteristic zero. For arbitrary $G$-bundles that may not admit a strongly canonical reduction, we have the following rough form of our another main result.
@@
The height filtration of $X$ is given by successively deleting some Frobenius twists of Schubert cells.
More explicitly, a theorem by Langer [7] shows for a principal $G$-bundle $F$ the strongly canonical reduction exists for $({\rm Fr}^{n})^{*}F$, when $n$ is sufficiently large. Therefore, consider the following Cartesian diagram for such $n$,
[…]
@@
$\tilde{X}=((\operatorname{Fr_{K}}^{n})^{*}F/P)_{K}$$X=(F/P)_{K}$${\rm Spec~}K$${\rm Spec~}K$$\scriptstyle{\phi_{K}}$$\scriptstyle{\operatorname{Fr}_{C}^{n}}$
where $\operatorname{Fr}_{C}$ is the absolute Frobenius on $C$.
@@
Suppose $n$ is large enough in the above sense, the height filtration of $h_{\mathcal{L}_{\lambda}}$ on $X$ is given by the image of the height filtration of the height filtration of $\tilde{X}$ along the homeomorphism $\phi_{K}$, the successive minima of $X$ are $\frac{1}{p^{n}}$ of the succesive minima of $\tilde{X}$.
@@
Here the height filtration of $\tilde{X}$ is known and computed by Theorem 3.1. Thus, we say the height filtration of $X$ is given by successively deleting some Frobenius twist of Schubert cells as stated in Theorem 1.4.
@@ 1.2. Toy example: projective spaces
Let $k$ be an algebraically closed field of positive characteristic and $C$ be a curve over $k$ with function field $K=k(C)$. Let $E$ be a vector bundle of rank $n$ on $C$. We take $\mathcal{X}=\mathbb{P}(E)$, $\mathcal{L}$ to be the relative ample line bundle $\mathcal{O}_{\mathbb{P}(E)}(1)$ and $X=\mathbb{P}(V)$ where $V$ is the generic fiber of $E$. A special case of a theorem by Ballaÿ [2] shows that: Let $\zeta_{ess}:=\inf_{t\in\mathbb{R}}\{Z_{t}=X\}$. Then
@@
where $\mu_{\max}(-)$ is the maximal slope appears in the Harder-Narasimhan filtration of $\pi_{*}\mathcal{L}^{\otimes n}$.
@@
In positive characteristic, the symmetric power of a semistable vector bundle may not remain semistable, making the left-hand side difficult to compute. This issue arises only in positive characteristic. However, this problem can be resolved by sufficiently twisting the Frobenius morphism multiple times.
@@ 1.3. Acknowledgments
We are grateful to Dr. Binggang Qu for his invaluable advice and support during the 2024 Algebra and Number Theory Summer School at Peking University. We also thank the organizers for bringing participants together and for their warm hospitality. The first author additionally thanks Prof. Jiu-Kang Yu for thoughtful comments. This work was largely written during the first author’s graduate study at The Chinese University of Hong Kong.
@@ 2.1. Vector bundles
Let $E$ be a vector bundle on a projective smooth curve $C$. The degree of $E$ is ${\mathrm{deg}}(E):={\mathrm{deg}}(\det(E))$ and the slope of $E$ is $\mu(E):={\mathrm{deg}}(E)/\operatorname{rk}(E)$. It is called slope semistable if for every subbundle $F$ of $E$, $\mu(F)\leq\mu(E)$. This is equivalent to $\mu(Q)\geq\mu(E)$ for every quotient bundle $Q$ of $E$.
There exists uniquely a filtration of subbundles $0=E_{0}\subseteq E_{1}\subseteq\cdots\subseteq E_{r}=E$ such that
@@
This filtration is called the Harder-Narasimhan filtration of $E$. $\mu(E_{1})$ is called maximal slope of $E$ and is also denoted by $\mu_{\max}(E)$. This theory can be generalized for any reductive group $G$, whereas in this case $G=GL_{n}$ by taking frame bundles of vector bundles of rank $n$.
@@ 2.2. Principal bundles
A principal $\Gamma$-bundle on $C$ is a $C$-scheme $F$ equipped with a right action of $\Gamma$ and a $\Gamma$-equivariant smooth morphism $F\longrightarrow C$ such that the map
@@
${\mathrm{deg}}(F):X(\Gamma)\longrightarrow\mathbb{Z},\quad\lambda\longmapsto\langle{\mathrm{deg}}(F),\lambda\rangle={\mathrm{deg}}(F\times_{\Gamma}k_{\lambda}),$
where ${\mathrm{deg}}(F\times_{\Gamma}k_{\lambda})$ is the degree of the line bundle $F\times_{\Gamma}k_{\lambda}$ on the curve $C$.
Let $H$ be a closed subgroup of a linear algebraic group $\Gamma$ over $k$. A reduction of structure group of $F$ to $H$ is a pair $(F_{H},\phi)$ where $F_{H}$ is a principal $H$-bundle and $\phi:F_{H}\times_{H}\Gamma\simeq F$ is an isomorphism. The quotient $F/H$ parametrizes all reductions of $F$ to $H$, see [1]. In particular, the assignment sending a section $\sigma:C\longrightarrow F/H$ to the reduction $\sigma^{*}F$ of $F$ to $H$ is a one-one correspondence between sections of $F/H\longrightarrow C$ and reductions of structure group of $F$ to $H$.
@@ 2.3. Reductive groups, characters and cocharacters
and by taking duals, we get the so-called slope map $X(P)^{\vee}_{\mathbb{Q}}\longrightarrow X(T)^{\vee}_{\mathbb{Q}}$ introduced in [10, §2.1.3].
2.4. The strongly canonical reduction
Let $G$ be a connected reductive group over $k$ and $F$ be a principal $G$-bundle over $C$. We denote $F(G):=F\times_{G,{\rm int}}G$ the group scheme associated to $F$ by the action of $G$ on itself by inner automorphisms. Let $F_{P}$ be a reduction of $F$ to a parabolic subgroup $P$. Let ${\mathrm{deg}}(F_{P})\in X(T)_{\mathbb{Q}}^{\vee}$ be the induced (rational) cocharacter. We also denote $F_{P}(P)=F_{P}\times_{P,{\rm int}}P$ the associated group scheme to $F_{P}$.
The $G$-bundle $F$ is called semistable if for any parabolic subgroup $P$, any reduction $F_{P}$ of $F$ to $P$ and any dominant character $\lambda$ of $P$ which is trivial on $Z(G)$, we have $\langle{\mathrm{deg}}(F_{P}),\lambda\rangle\leq 0$, or equivalently, if any parabolic subgroup scheme of $F(G)$ has non positive-degree.
@@ 2.4. The strongly canonical reduction
Remark 2.2.
Semistability is preserved by separable pullbacks, that is, for a separable morphism $f:C^{\prime}\to C$, the pullback $f^{*}(F_{P})$ is semistable. However, this fact does not hold for inseparable morphism.
Among all filtrations of a vector bundle, there is a canonical one (the Harder-Narasimhan filtration). More generally, among all reduction of a principal $G$-bundles to parabolic subgroups, there is also a canonical one. The reduction $E_{P}$ of $E$ is called canonical (or the Harder–Narasimhan filtration) if it satisfies the following conditions:
A reduction $F_{P}$ of $F$ to a parabolic subgroup $P$ is called canonical if the following two conditions hold:
@@
We say $F_{P}$ is strongly canonical if for any non-constant finite morphism $f:C^{\prime}\to C$, the pullback $f^{*}(F_{P})$ is canonical.
@@
Similarly as strongly Semistability, the notion of canonical reduction is preserved by separable pullbacks [3]. In characteristic zero, canonical reduction is strong while this does not hold for inseparable morphism, which makes Definition 2.3 necessary.
We will need the following theorem.
Theorem 2.5 ([8], Theorem 3.23).
@@ 3. Height filtrations and successive minima
Let $F_{Q}$ be the canonical reduction of $F$ to $Q$, and ${\mathrm{deg}}(F_{Q})\in X(T)^{\vee}_{\mathbb{Q}}$ be the degree cocharacter. Let $W,W_{P}$ and $W_{Q}$ be the Weyl group of $G,L_{P}$ and $L_{Q}$. For $w\in W_{Q}\backslash W/W_{P}$, let $C_{w}=(F_{Q}\times_{Q}QwP/P)_{K}\subseteq X$ be the corresponding Schubert cell in $X$. We have the following theorem.
@@
In particular, the successive minima are $\zeta_{w}=\left\langle\operatorname{deg}\left(F_{Q}\right),w\lambda\right\rangle$ and Zhang’s successive minima are $e_{i}=\min\left\{\zeta_{w}:\ell(w)=\operatorname{dim}G/P-i+1\right\}$ where
@@
$\tilde{X}=((\operatorname{Fr_{K}}^{n})^{*}F/P)_{K}$$X=(F/P)_{K}$${\rm Spec~}K$${\rm Spec~}K$$\scriptstyle{\phi_{K}}$$\scriptstyle{\operatorname{Fr}_{C}^{n}}$